OpenStudy (anonymous):

A rope is being used to anchor a tree to the ground. If the rope is 15 feet long and it is tied to the tree 12 feet above the ground, how far from the base of the tree is the rope tied to the ground? 19 ft 3 ft 6 ft 9 ft

OpenStudy (anonymous):

please i realy need helpl!!!!

OpenStudy (anonymous):

pythagoras for this one

OpenStudy (anonymous):



OpenStudy (anonymous):

\[x^2+12^2=15^2\] \[x^2=15^2-12^2\] \[x^2=225-144\] \[x^2=81\] \[x=9\]

OpenStudy (anonymous):

with a little practice, you will recognize this as the famous 3 - 4 - 5 right triangle in disguise. instead of 3 - 4 - 5 if you multiply all by 3 you get 9 - 12 - 15

OpenStudy (anonymous):

so it is 9ft? thanks!!! much

OpenStudy (anonymous):

in other words after a while when you see for example a hypotenuse of 25 and a side as 20 you will know right away that the other side is 15 because it is still the 3-4-5 right triangle with everything multiplied by 5

OpenStudy (anonymous):

yw